OpenStudy (vishweshshrimali5):

V.I. Arnold's Problem: From a barrel of wine, a spoon was poured into a cup of tea, and then the same spoon of the obtained (non homogeneous!) mixture was poured from the cup back into the barrel. Where did the amount of the foreign beverage become greater ?

OpenStudy (vishweshshrimali5):

@ganeshie8 @Hero @hartnn @Kainui @iambatman @mathslover

OpenStudy (inkyvoyd):

how does one pour a spoon? do you mean a spoonful?

OpenStudy (vishweshshrimali5):

Well I think he meant he just first inserted the spoon in one container and then the other one and finally again in the first one.

OpenStudy (vishweshshrimali5):

I think that this must depend on the properties of the liquid. The larger the amount of the liquid particles stuck to the spoon, the higher must be the amount of foreign beverage.

OpenStudy (vishweshshrimali5):

But, this must make the problem a chemistry or a physics problem and its a maths question and he says that "children 4-5 years old would have solved them in half an hour"

OpenStudy (vishweshshrimali5):

@Miracrown

ganeshie8 (ganeshie8):

One full spoon of `pure wine` was poured into the glass One full spoon of `tea+wine` mixture was poured back into the barrel

ganeshie8 (ganeshie8):

hmm some wine that was poured earlier into the glass was taken back into the barrel as well

OpenStudy (vishweshshrimali5):

Yeah if he meant spoonful then its easy to solve. But, what did he mean !!!

ganeshie8 (ganeshie8):

spoonful makes more sense...

OpenStudy (vishweshshrimali5):

Yeah. I agree with you @ganeshie8 . If that's case, then, cup of tea must be the answer.

ganeshie8 (ganeshie8):

maybe, we need to prove it

OpenStudy (anonymous):

He's asking in which (cup or barrel) is the foreign liquid greater in.

OpenStudy (vishweshshrimali5):

May be. But, I think I can do that.

OpenStudy (vishweshshrimali5):

Let, the volume of the liquid in one spoonful = V mL Then, the wine transferred in cup of tea due to first step = V mL. Let, a cup of tea had tea = x mL and wine present initially = y mL Then, cup has x mL tea + V mL wine after first step Now, second step will transfer V mL mixture into barrel of wine. Now, since V < x, thus, amount of wine would be less than amount of tea in the spoon. But, even then the amount of tea would be less than V. So, after second step the barrel of wine would contain: (y - V + wine from spoon) mL wine + (tea from spoon and this amount < V) Clearly, foreign amount in tea cup = V mL foreign amount in barrel of wine < V Thus, tea of cup would have more foreign amount

OpenStudy (vishweshshrimali5):

What do you think @ganeshie8 ?

OpenStudy (vishweshshrimali5):

Here I would love to share this link for young students: http://jnsilva.ludicum.org/HMR13_14/Arnold_en.pdf (I recently found it after a lot of research)

OpenStudy (anonymous):

This is probably just a tricky question, where the answer is they both have equal amounts of the foreign substance.

OpenStudy (vishweshshrimali5):

@iambatman But, I don't think they can have the same amount. Even if you think in terms of probability, then also the amount of wine in the spoon would be very less after second step but not zero.

ganeshie8 (ganeshie8):

key thing is to notice that the volumes of barrel and glass don't change after the experiment

OpenStudy (anonymous):

That's what I'm thinking, if you think of its initial state, there's no change.

OpenStudy (vishweshshrimali5):

Yeah that's correct.

ganeshie8 (ganeshie8):

volume of water entered into the barrel = volume of wine left in the glass

ganeshie8 (ganeshie8):

so i think iambatman has it !

OpenStudy (vishweshshrimali5):

But, volume of the liquid is constant not the type of liquid. Isn't it ?

OpenStudy (nincompoop):

ASK me

OpenStudy (vishweshshrimali5):

Ok @nincompoop ?

OpenStudy (vishweshshrimali5):

Well I think I got it. Thanks @iambatman and @ganeshie8 .

OpenStudy (nincompoop):

first, post this in Chemistry section

OpenStudy (nincompoop):

then I will break it down

OpenStudy (vishweshshrimali5):

Why ?

OpenStudy (anonymous):

Np, glad I could help, we often tend to over complicate the problems, such as nins chess problem, yeah I said it ;).

OpenStudy (nincompoop):

because you used the word non-homogenous mixture

ganeshie8 (ganeshie8):

chess is complicated

OpenStudy (kainui):

pretend the mixture is a differential equation, and solve it here.

OpenStudy (anonymous):

Ik ganeshie, I was joking :p

mathslover (mathslover):

Oh I missed this one.. :( Sorry, I couldn't help you this time @vishweshshrimali5 :P lol

OpenStudy (vishweshshrimali5):

That's it @Kainui. I was thinking of finding something to infuriate my cousins by asking them this question and saying that its very difficult; you have to use DE. :D

OpenStudy (nincompoop):

you use chemistry tricks

OpenStudy (vishweshshrimali5):

Number of moles transferred by assuming molarity ?

OpenStudy (nincompoop):

concentration and dilution

OpenStudy (vishweshshrimali5):

Ahh M*V = constant !! Well, this could also help. Thanks @nincompoop .

OpenStudy (nincompoop):

the units you will use are barrel, cup and spoon

OpenStudy (nincompoop):

use this hypothetical conversion units \[1Barrel_w \times \frac{ 1000 cupful_w}{ 1Barrel_w } \times \frac{ 50 spoonful_w }{ 1 cupful_w }=50000spoonful_w\] and since you're familiar with chem already, then compare the amounts just like how mole fractions are done \[MolefractionA= \chi_A = \frac{ n_A }{ n_A+n_B }\]

OpenStudy (vishweshshrimali5):

OKay I will try solving the question using these units. Thanks again @nincompoop